# Finding Dori [Uncaptioned image]: Memorization in Text-to-Image Diffusion Models Is Less Local Than Assumed

### v1

@@ G.1 Locality in the text embedding space
G.2 Locality in the model’s weights
@@
$\text{Discrepancy}({\bm{y}}_{i},{\bm{y}}_{j})=||\text{Activations}({\bm{y}}_{i})-\text{Activations}({\bm{y}}_{j})||^{2}_{2}\text{,}$
where Activations$({\bm{y}})$ outputs a vector of activations of a given layer further used to identify memorization weights by Wanda or NeMo. For NeMo, we obtain activations from passing the text embedding ${\bm{y}}$ through the value layer in cross-attention blocks, and Wanda utilizes activation of the feed-forward layer after the attention operator in cross-attention blocks. To assess mean pairwise discrepancy of in set $\bm{Y}=\{{\bm{y}}_{i}|i=1,\dots,N\}$ of size $N$ we use
@@
where ${\bm{y}}_{i}$ and ${\bm{y}}_{j}$ are two embeddings that trigger replication of some memorized image(s), and Weights$({\bm{y}})$ returns a set of weights identified by a pruning-based mitigation method, #$\#\bm{Y}$ denotes the size of the set. Analogically to MeanDiscrepancy, we define the mean pairwise weight agreement for a set of embeddings $\bm{Y}$ as
@@
G.3 NeMo and Wanda identify memorization weights in different layers
We examine the behavior of pruning-based methods through the lens of their weights selection. To this end, we compute these weights for all VM samples, separately for each memorized image. In Fig. 8 we show that NeMo tend to identify memorization weights only in four out of seven layers. This result explains high weight agreement in layers two, six, and seven in Fig. 3 (right), since when no weights are identified in a layer, we set agreement to 1. Results for Wanda contrast with the results for NeMo, as it finds more traces of memorized content in deeper layers of the model (five, six, and seven). Importantly, in these layers also the agreement drops significantly, as can be seen in Fig. 3 (right).
Figure 8: Number of memorization weights per layer. We observe that for NeMo, no weights are identified to prune in layers two, six, and seven (left). Conversely, Wanda identifies significantly more memorization weights in deeper layers. Interestingly, the drop in weight agreement for Wanda (Fig. 3) happens also in the deeper layers of the model.
Appendix H Qualitative Results
H.1 Qualitative Results for Wanda Pruning
@@
H.2 Qualitative Results for NeMo Pruning
@@
H.3 Adversarial Fine-Tuning
@@
H.4 Wanda with 10% Sparsity
@@
H.5 Qualitative Results for Iterative NeMo Pruning

### v2

@@ I.7 Our Mitigation Does not Cause Unintended Memorization of Other Samples
Additionally, we compute pairwise L2 distance in the embedding space for (1) random initialization ($\mathcal{N}(\bm{0},\bm{I})$), (2) adversarial embeddings optimized from $\mathcal{N}(\bm{0},\bm{I})$, (3) embeddings of non-memorized prompts (${\bm{y}}_{nonmem}$), (4) adversarial embeddings optimized from ${\bm{y}}_{nonmem}$. To our surprise, the adversarial embeddings that trigger generation of the same memorized samples appear to be more spread out than randomly initialized embeddings, and are also more spread out than embeddings of non-memorized prompts, as it is visible in Fig.˜9 (left).
Interestingly, initialization of optimization from $\mathcal{N}(\bm{0},\bm{I})$ is more beneficial to finding Dori. We observe that the embeddings have to be changed less than when we initialize them from prompts, which is expressed by lower L2 distance between initializations and the final embeddings ${\bm{y}}_{adv}$ in Fig.˜9 (right).
Figure 9: L2 distances of input embeddings. Left: we compute pairwise L2 distances in text embedding space within the set of $100$ random embeddings ($\mathcal{N}(\bm{0},\bm{I})$), set of adversarial embeddings optimized from random embeddings (second box from the left), $100$ randomly selected non-memorized prompts (${\bm{y}}_{nonmem}$) and adversarial embeddings optimized from non-memorized embeddings (fourth box from the left). We observe that after optimization, the adversarial embeddings are more spread out in the text embedding space than their initial points (be it $\mathcal{N}(\bm{0},\bm{I})$ or randomly selected ${\bm{y}}_{nonmem}$). Right: We compute the L2 distance between the initialization and the final adversarial embeddings. We note that when initializing the optimization with ${\bm{y}}_{nonmem}$ we have to travel farther in the text embedding space to obtain an adversarial embedding ${\bm{y}}_{adv}$ that successfully triggers replication of the memorized image ${\bm{x}}_{\mathit{mem}}$.
J.2 Locality in the model’s weights
@@
[…]
@@
[…]
@@
J.3 Locality is not Present in Any Layers
Our findings from Section˜4.3 suggest that locality of memorized data in the specific layers of the model is an illusion. In this section, we extend the analysis to other layers, specifically Q, K, V, and MLP layers in self- and cross-attention modules, as well as convolutional layers in ResNet modules of the U-Net. To evaluate locality, we default to activations, specifically the MeanDiscrepancy, defined in Section˜J.2, and omit weight agreement, since neither Wanda nor NeMo focuses their mitigation efforts on these additional layers.
The results in Fig.˜10 further undermine the notion of locality. In all layer types, we observe high discrepancy for various adversarial embeddings ${\bm{y}}_{adv}$ associated with the same image. This shows that trying to localize neurons for pruning memorization mitigation based on activations can not work, as vastly different activation patterns lead to the same memorized image.
Interestingly, the discrepancy for ${\bm{y}}_{\mathit{mem}}$ associated with different images is significantly higher for layers that interact with the image features directly (like Q matrices in self-attention), and the effect strengthens deeper in the model. We argue that this is an expected behavior—outputs (and thus activations) of the model should diverge for different images.
Figure 10: Locality assumption fails in all layers of the U-Net. We observe high activation discrepancy for adversarial embeddings triggering the same image, regardless of the layer type and depth.
J.4 NeMo and Wanda identify memorization weights in different layers
We examine the behavior of pruning-based methods through the lens of their weights selection. To this end, we compute these weights for all VM samples, separately for each memorized image. In Fig.˜11 we show that NeMo tend to identify memorization weights only in four out of seven layers. This result explains high weight agreement in layers two, six, and seven in Fig.˜4, since when no weights are identified in a layer, we set agreement to 1. Results for Wanda contrast with the results for NeMo, as it finds more traces of memorized content in deeper layers of the model (five, six, and seven). Importantly, in these layers also the agreement drops significantly, as can be seen in Fig.˜4.
Figure 11: Number of memorization weights per layer. We observe that for NeMo, no weights are identified to prune in layers two, six, and seven (left). Conversely, Wanda identifies significantly more memorization weights in deeper layers. Interestingly, the drop in weight agreement for Wanda (Fig.˜4) happens also in the deeper layers of the model.
Appendix K Qualitative Results
K.1 Qualitative Results for Wanda Pruning
@@
K.2 Qualitative Results for NeMo Pruning
@@
K.3 Adversarial Fine-Tuning
@@
K.4 Wanda with 10% Sparsity
@@
K.5 Qualitative Results for Iterative NeMo Pruning